OpenStudy (anonymous):

Find the derivative of ln(8x)

OpenStudy (anonymous):

1/x

OpenStudy (anonymous):

An easy way to see that is to remember that: \[\ln (8x) = \ln(8) + \ln(x)\] Then, because "ln(8)" is just a constant, it disappears when you take the derivative...